OpenStudy (anonymous):

find by implicit differentiation dy/dx : (3e^xy) - x = 0

OpenStudy (raden):

is it like this : 3e^(xy) - x = 0 ? if yes, differentiation implicitly is 3ye^(xy) + 3xe^(xy) y ' - 1 = 0 3xe^(xy) y ' = 1 - 3ye^(xy) y ' = {1 - 3ye^(xy)}/3xe^(xy)